OpenStudy (anonymous):

Find the derivative of the function. f(x) = (8e^x)-(x^2e^x)

OpenStudy (anonymous):

Hint (1) : derivative of e^x = e^x. Hint (2) : Deriving multiplied values: The first times the derivative of the second plus the second times the derivative of the first. Learn that phrase.

OpenStudy (anonymous):

ok so the derivative 8e^x-2e^x?

OpenStudy (anonymous):

f'(x)=d/dx(8e^x)-d/dx(x^2e^x) you will need to use the product rule for the second one 8e^x-(d/dx(x^2)e^x+x^2(d/dx(e^x))) 8e^x -(2xe^x+x^2e^x) 8e^x-2xe^x-x^2e^x e^x(8-2x-x^2)